OpenStudy (anonymous):

Determine the point of intersection between the lines (x-3,y-20,z-7)=t(2,-4,5) ter, and (x-5)/2 = y-2 = (z+4)/-3 - AMISTRE64 HELP PLEASE!

OpenStudy (anonymous):

answer is no intersection / skew lines but how?!

OpenStudy (anonymous):

yay amistre :D !

OpenStudy (amistre64):

since they are in 3d space; they can look to cross, but actually never touch. such as the line in the ceiling and the line on the floor both make it to a corner but never intersect each other

OpenStudy (anonymous):

how do you prove that mathematically. i do understand that concept that you just mentioned

OpenStudy (anonymous):

i can get the second equation in terms of x y and z but how do i sub that in to the second equation and how do i deal with t(2,-4,5) ? !

OpenStudy (anonymous):

that is how far i have went ... just stuck at that step

OpenStudy (amistre64):

we prove it by seeing if the scalar for one line vector equates to the scalar in the other ... but i gotta review it real quick to recall some details

OpenStudy (anonymous):

kay take your time boss :)

OpenStudy (amistre64):

when lines intersect... (x-3,y-20,z-7) = t(2,-4,5) ter ?ter? whats this?

OpenStudy (anonymous):

its one of the two equationse that intersect .. the format of this eqaution has confused memyself

OpenStudy (amistre64):

whats the ?ter? mean i take it its a typo of some kind

OpenStudy (anonymous):

ter means where value of t can me any real number sorry .. u knw like xer

OpenStudy (amistre64):

(x-5)/2 = y-2 = (z+4)/-3 a=2; b=1 ; c=-3 ; just the denoms of your things here right?

OpenStudy (amistre64):

t<2,1,-3> is the vector and P(5,2,-4)

OpenStudy (anonymous):

yeah denoms are right

OpenStudy (amistre64):

x = 5 +2t y = 2 + t z = -3 -4t for our equation of a line in R^3 or at least one of them

OpenStudy (anonymous):

you mean z = -4-3t ?

OpenStudy (amistre64):

(x-3,y-20,z-7)=t(2,-4,5) x= 3 +2t y= 20 -4t z= 7 +5t is our other line right?

OpenStudy (amistre64):

yeah :) lol

OpenStudy (anonymous):

hmm but how did you get that parametric equation ... you just diregarded x- y- etc?

OpenStudy (amistre64):

now we equate our x=x but use a different variable to indicate our scalar; say s instead of t and solve for the system of equations

OpenStudy (amistre64):

x-3 means x = 3 + vector x component

OpenStudy (amistre64):

x-x0 = Vx x = x0 + Vx

OpenStudy (anonymous):

oh true indeed

OpenStudy (amistre64):

5 +2t = 3 +2s 2 + t = 20 -4s -4 -3t = 7+5s

OpenStudy (amistre64):

solve for s and t and see if they match in the 3rd spot

OpenStudy (anonymous):

hmm :O i can work the rest you can help other people as with with great power comes great responsibilities :P

OpenStudy (amistre64):

if i start sticking to the walls ........

OpenStudy (amistre64):

it prolly means i need a bath lol

OpenStudy (anonymous):

lmfao xD

OpenStudy (amistre64):

if the system of equations creates s and t to meet at a common point; then they intersect there; if not then they never meet

OpenStudy (anonymous):

GOT IT YAY :) ! Love you man .. studying to get int queens university bro :(

OpenStudy (anonymous):

i got it just got to maintain their conditional offer :)